OpenStudy (anonymous):

find the solution of this system of equations. Separate the x- and y- values with a comma. x=2+y 8x-2y=22

OpenStudy (anonymous):

@shortycme

OpenStudy (anonymous):

We can plug in '2 + y' for 'x' in the 2nd equation: 8x - 2y = 22 8(2 + y) - 2y = 22 Distribute 8 into the parenthesis: 10 + 8y - 2y = 22 Subtract 8y - 2y @noelle34

OpenStudy (anonymous):

answer?

OpenStudy (anonymous):

I am not going to give the answer, you must participate if you want the answer.

OpenStudy (anonymous):

then bye

syed98 (syedmohammed98):



syed98 (syedmohammed98):

I hope you get it :)

OpenStudy (anonymous):

If I got it I wouldn't be asking in the first place.

syed98 (syedmohammed98):

well okay....just trying to help....:)